OpenStudy (anonymous):

Write the point-slope equation of the line through (-2, 1) with a slope of 3. A.y - 1 = 3(x + 2) B.y - 2 = 3(x + 1) C.y + 2 = 3(x - 1)

OpenStudy (anonymous):

plug in the value of slope in y=mx+c where m is slope and c is to be found by plugging value of x and y that is (-2,1)

OpenStudy (anonymous):

point slope is \[y-y_1=m(x-x_1)\] in your example \(x_1=-2,y_1=1, m=3\)

OpenStudy (anonymous):

ok so... y-1=3(x+2)

OpenStudy (anonymous):

yes simplify it to get your answer

OpenStudy (anonymous):

no no, do not "simplify" anything make a direct substitution

OpenStudy (anonymous):

wait what? Isn't the answer A?

OpenStudy (anonymous):

yes A is the answer

OpenStudy (anonymous):

\[\large y-\color {red}{y_1}=\color{blue}m(x-\color{green}x_1)\]

OpenStudy (anonymous):

Ok. Thank you guys for your help!